# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 2.1. Some basic knowledge
([7], Theorem 3.2.11) Let $P$ be a geodesic in $\mathbb{H}^{2}$ and let $y\in\mathbb{H}^{2}$ be a time-like vector. Then there exists a unique geodesic of $\mathbb{H}^{2}$ through $y$ and Lorentz orthogonal to $P$.
2.2. A geometric explanation
Consider a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ in the Kelin model. Set $\{r,s,t\}=\{i,j,k\}$. Set $E_{rs}=\mathrm{Span}(v_{r},v_{s})\cap\mathbb{H}^{2}$, then $E_{rs}$ is a geodesic in $\mathbb{H}^{2}$. Set $v_{r}^{\bot}\cap\mathbb{H}^{2}=E^{\prime}_{st}$, then $E^{\prime}_{st}$ is also a geodesic in $\mathbb{H}^{2}$ and Lorentz orthogonal to the geodesics $E_{rs},E_{rt}$. The geodesic arc in $E^{\prime}_{st}$ bounded by $E_{rs}$ and $E_{rt}$ is called a boundary arc, denoted by $r\in B$. The geodesic arc between two adjacent boundary arcs $r$ and $s$ is denoted by $\{rs\}$, which is a part of $E_{rs}$ and called an ideal edge. Moreover, the six geodesic arcs form a right-angled hyperbolic hexagon $\{ijk\}$. Conversely, for any geodesic $E_{rs}$ in $\mathbb{H}^{2}$, there exists a unique space-like vector $v^{\prime}_{t}$ such that $(v^{\prime}_{t})^{\bot}\cap\mathbb{H}^{2}=E_{rs}$. Then the three points $v^{\prime}_{i},v^{\prime}_{j}$ and $v^{\prime}_{k}$ are hyper-ideal and form a hyper-ideal hyperbolic triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, which is called the polar triangle of the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$. Please refer to Figure 2.
Figure 2. A hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and its polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ in the Kelin model (left), and a right-angled hyperbolic hexagon $\{ijk\}\in F$ in the Poincaré model (right).
By the construction of the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, the vertices $v^{\prime}_{i},v^{\prime}_{j},v^{\prime}_{k}$ are determined by the vertices $v_{i},v_{j},v_{k}$ and vice versa. Without loss of generality, we assume $||v_{i}||=||v_{j}||=||v_{k}||=1$, then
(12) $v^{\prime}_{i}=\frac{v_{k}\otimes v_{j}}{||v_{k}\otimes v_{j}||},\ v^{\prime}_{j}=\frac{v_{i}\otimes v_{k}}{||v_{i}\otimes v_{k}||},\ v^{\prime}_{k}=\frac{v_{j}\otimes v_{i}}{||v_{j}\otimes v_{i}||}.$
Therefore, given a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, there exists a unique polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ with its vertices satisfying (12). The length of the ideal edge $\{rs\}\in E$ is denoted by $l_{rs}$ and the length of the boundary arc $t\in B$ is denoted by $\theta^{rs}_{t}$.
@@ 2.2. A geometric explanation
The vectors $c_{ijk},c_{rs},v^{\prime}_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$, where the notations of $c_{ijk},c_{rs}$ follows from Remark 1.4.
@@
Without loss of generality, we assume $r=i,s=j,t=k$. Since $c_{ij}\in E_{ij}$, by Proposition 2.4, there exists a unique geodesic of $\mathbb{H}^{2}$ through $c_{ij}$ and Lorentz orthogonal to $E_{ij}$, which is $P_{ij}=\mathrm{Span}(c_{ij},c_{ijk})\cap\mathbb{H}^{2}$. Note that $E_{ij}=(v^{\prime}_{k})^{\bot}\cap\mathbb{H}^{2}$. By Proposition 2.3, since $E_{ij}$ and $P_{ij}$ meet at a right angle, then $v^{\prime}_{k}\in\mathrm{Span}(c_{ij},c_{ijk})$. Q.E.D.
Combining Remark 1.4 and Lemma 2.5, we have $c_{ijk}=\mathrm{Span}(v^{\prime}_{i},c_{jk})\cap\mathrm{Span}(v^{\prime}_{j},c_{ik})\cap\mathrm{Span}(v^{\prime}_{k},c_{ij})$. Due to the equivalence of $v_{i}v_{j}v_{k}$ and $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it is natural to consider a dual model. Note that for any $c_{ijk}\in\mathbb{R}^{3}$, $\mathrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}$ may not intersect within $\mathbb{H}^{2}$. For the convenience of calculations, we focus on the case that
$\mathrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}\cap\mathbb{H}^{2}=\mathrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}\neq\emptyset,\quad\forall r\in\{i,j,k\}.$
This condition limits the positions of $c_{ijk}$ when $c_{ijk}$ is space-like, i.e., $c_{ijk}$ can only be in some certain open domains. Denote $c^{\prime}_{st}=\mathrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}$ as the dual edge center, which is the unique point along the geodesic $E^{\prime}_{st}$ such that $c^{\prime}_{st}$ is of (signed) distance $\theta_{st}$ from $\{rt\}$ and $\theta_{ts}$ from $\{rs\}$. Thus we have
@@
Note that the signed distance $\theta_{st}$ is positive if $c^{\prime}_{st}$ is on the same side of $\{rt\}$ as $\{rs\}$ along the hyperbolic geodesic $E^{\prime}_{st}$, and negative otherwise. The above construction shows the following lemma.
@@
The vectors $c_{ijk},c^{\prime}_{rs},v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
By Proposition 2.4, the edge perpendicular $P^{\prime}_{ij}=\mathrm{Span}(v_{k},c_{ijk})\cap\mathbb{H}^{2}$ is through $c^{\prime}_{ij}$ and Lorentz orthogonal to $E^{\prime}_{ij}$. Hence, the three edge perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ also intersect in a point $c_{ijk}$. Similar to [16, Lemma 2.4], we further have that the three perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ intersect in a point if and only if
(14) $\sinh\theta_{ij}\sinh\theta_{jk}\sinh\theta_{ki}=\sinh\theta_{ji}\sinh\theta_{kj}\sinh\theta_{ik}.$
Hence, Definition 1.1 is equivalent to the following definition with respect to the parameter $\theta$. For precision, we rewrite the boundary component $i\in B$ as $\{j^{\prime}k^{\prime}\}\in B$. We use $B_{+}$ to denote the sets of oriented boundary components and label them with ordered pairs $(j^{\prime},k^{\prime})\in B_{+}$.
Definition 2.7.
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. An assignment of partial edge lengths is a map $\theta:B_{+}\rightarrow\mathbb{R}$, such that $\theta_{i}^{jk}=\theta_{jk}+\theta_{kj}>0$ for each oriented boundary arc $(j^{\prime},k^{\prime})\in B_{+}$ and formula (14) holds for every right-angled hyperbolic hexagon $\{ijk\}\in F$.
Recall the definition of $h_{rs}$, which is the signed distance of the face center $c_{ijk}$ to the hyperbolic geodesic $E_{rs}$ and is positive if $c_{ijk}$ is on the same side of the geodesic $E_{rs}$ as the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E_{ij}$). Similarly, we denote $q_{rs}$ as the signed distance from $c_{ijk}$ to $c^{\prime}_{rs}$, which is defined to be positive if $c_{ijk}$ is on the same side of the geodesic $E^{\prime}_{rs}$ as the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E^{\prime}_{rs}$). Please refer to Figure 3.
@@
In the Klein model, the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ divide the unit disk into 13 open domains, denoted by $D_{1},D_{2},...,D_{13}$ respectively. Please refer to Figure 4.
Figure 4. Division for a right-angled hyperbolic hexagon $\{ijk\}$ in the Klein model. If $c_{ijk}$ is space-like, then it can be in red and blue open domains.
By the definition of $h$ and $q$, one can directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 13 open domains, as shown in Table 1. For example, suppose $c_{ijk}$ is in the domain $D_{1}$. Then $c_{ijk}$ is on the opposite side of $E_{ik}$ as $v_{i}v_{j}v_{k}$ and hence $h_{ik}<0$. And $c_{ijk}$ is on the same side of $E_{ij},E_{jk}$ as $v_{i}v_{j}v_{k}$ and hence $h_{ij}>0,h_{jk}>0$. Similarly, $c_{ijk}$ is also on the opposite side of $E^{\prime}_{ij}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and hence $q_{ij}<0$. And $c_{ijk}$ is on the same side of $E^{\prime}_{ik},E^{\prime}_{jk}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and hence $q_{jk}>0,q_{ik}>0$. If $c_{ijk}$ is space-like, then $c_{ijk}$ can be only in one of 12 open domains and each domain is bounded by two tangential lines and unit circle, denoted by $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ and $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ respectively. Please refer to Figure 4. If $c_{ijk}$ is not in one of 12 open domains, then either $\mathrm{Span}(c_{ijk},v_{r})\cap\mathbb{H}^{2}=\emptyset$ or $\mathrm{Span}(c_{ijk},v^{\prime}_{r})\cap\mathbb{H}^{2}=\emptyset$, which contradicts with the assumption that the edge centers $c_{st}\in E_{rs}$ and $c^{\prime}_{st}\in E^{\prime}_{rs}$ are in $\mathbb{H}^{2}$ for $\{r,s,t\}=\{i,j,k\}$. Similar to the case that $c_{ijk}$ is time-like, one can also directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 12 open domains, as shown in Table 1. By moving $c_{ijk}$ to the light cone continuously, one can obtain similar results. Note that $c_{ijk}$ can only be on the unit disk that removes 12 points.
@@
In the following, we provide another way to obtain Table 1. Suppose $c_{ijk}$ is time-like. Without loss of generality, we assume $c_{ijk}$ is in the domain $D_{1}$. By Lemma 2.5, the vectors $c_{ijk},c_{ik},v^{\prime}_{j}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ik}\in E_{ik}$ is not in $\{ik\}$. Furthermore, since $c_{ijk}$ is in $D_{1}$, then $c_{ik}$ is on the opposite side of $k$ as $i$. By the definition of $d$ in Remark 1.3, we have $d_{ki}<0,d_{ik}>0$. Similarly, we have $d_{kj}<0,d_{jk}>0,d_{ij}>0,d_{ji}>0$. By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{ij},v_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, then $c^{\prime}_{ij}\in E^{\prime}_{ij}$ is not in $k$. Furthermore, since $c_{ijk}$ is in $D_{1}$, then $c^{\prime}_{ij}$ is on the opposite side of $\{ik\}$ as $\{jk\}$ and hence $\theta_{ji}<0,\theta_{ij}>0$. Similarly, we have $\theta_{jk}<0,\theta_{kj}>0,\theta_{ik}>0,\theta_{ki}>0$. By Corollary 2.11, we have $h_{ik}<0,h_{jk}>0,h_{ij}>0$ and $q_{ij}<0,q_{ik}>0,q_{jk}>0$. The case that $c_{ijk}$ is space-like follows from the following two lemmas.
@@
The signs of $q$ and $h$ in the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ are the same as those in the domains $D_{2},D_{4},D_{6},D_{8},D_{10},D_{12}$, respectively.
@@
We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{V}}$ are the same as those in the domain $D_{10}$, and the remaining cases are similar.
If $c_{ijk}$ is in the domain $D_{10}$, then $q_{jk}<0$ and the rest is positive by Table 1. Suppose $c_{ijk}$ is in the domain $D_{\mathrm{V}}$. By Lemma 2.5, the vectors $c_{ijk},c_{ij},v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ij}\in E_{ij}$ is not in $\{ij\}$. Furthermore, since $c_{ijk}$ is in $D_{\mathrm{V}}$, then $c_{ij}$ is on the opposite side of $j$ as $i$ and hence $d_{ij}<0,d_{ji}>0$. Similarly, we have $d_{ik}<0,d_{ki}>0,d_{jk}>0,d_{kj}>0$. By Proposition 2.12, we have $h_{ij}>0,h_{jk}>0,h_{ik}>0$. Note that $c_{ijk}$ is in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$. By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{rs},v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$ for $\{r,s,t\}=\{i,j,k\}$. Then $c^{\prime}_{rs}\in E^{\prime}_{rs}$ is in $t$. Hence, $\theta_{ij}>0,\theta_{ji}>0,\theta_{jk}<0,\theta_{kj}>0,\theta_{ik}>0,\theta_{ki}>0$. By Proposition 2.12, we have $q_{ij}>0,q_{jk}<0,q_{ik}>0$. Q.E.D.
@@
The signs of $q$ and $h$ in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ are the same as those in the domains $D_{1},D_{3},D_{5},D_{7},D_{9},D_{11}$, respectively.

### v3

@@ 2.1. Some fundamental results
([8], Proposition 4) Let $x\in\mathbb{R}^{3}$ be a space-like vector, and let $\gamma=x^{\bot}\cap\mathbb{H}^{2}$ be the geodesic corresponding to $x^{\bot}$. Suppose $\omega$ is another geodesic intersecting $\gamma$, and $P$ is the plane such that $\omega=P\cap\mathbb{H}^{2}$. Then $\gamma$ and $\omega$ intersect at a right angle if and only if $x\in P$.
2.2. A geometric interpretation
Let $v_{i}v_{j}v_{k}$ be a hyper-ideal hyperbolic triangle in the Klein model, and let $\{r,s,t\}=\{i,j,k\}$. By the notations in Remark 1.3, $E_{rs}=\textrm{Span}(v_{r},v_{s})\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$. Define $E^{\prime}_{st}=v_{r}^{\bot}\cap\mathbb{H}^{2}$. Then $E^{\prime}_{st}$ is also a geodesic in $\mathbb{H}^{2}$, and is Lorentz orthogonal to both $E_{rs}$ and $E_{rt}$. The geodesic arc in $E^{\prime}_{st}$ bounded by $E_{rs}$ and $E_{rt}$ is a boundary arc, denoted by $r$. An ideal edge $\{rs\}$ is the geodesic arc between two adjacent boundary arcs $r$ and $s$, lying on $E_{rs}$. Furthermore, the six geodesic arcs form a right-angled hyperbolic hexagon $\{ijk\}$. Conversely, for any geodesic $E_{rs}$ in $\mathbb{H}^{2}$, there exists a unique space-like vector $v^{\prime}_{t}$ such that $(v^{\prime}_{t})^{\bot}\cap\mathbb{H}^{2}=E_{rs}$. The three points $v^{\prime}_{i},v^{\prime}_{j},v^{\prime}_{k}$ are hyper-ideal and form a hyper-ideal hyperbolic triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, referred to as the polar triangle of $v_{i}v_{j}v_{k}$. See Figure 2 for illustration.
Figure 2. A hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and its polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ in the Klein model (left), and a right-angled hyperbolic hexagon $\{ijk\}\in F$ in the Poincaré disk model (right).
By the construction of the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, the vertices $v^{\prime}_{i},v^{\prime}_{j},v^{\prime}_{k}$ are determined by the vertices $v_{i},v_{j},v_{k}$, and vice versa. Without loss of generality, we assume $\|v_{i}\|=\|v_{j}\|=\|v_{k}\|=1$. Then
(11) $v^{\prime}_{i}=\frac{v_{k}\otimes v_{j}}{\|v_{k}\otimes v_{j}\|},\ v^{\prime}_{j}=\frac{v_{i}\otimes v_{k}}{\|v_{i}\otimes v_{k}\|},\ v^{\prime}_{k}=\frac{v_{j}\otimes v_{i}}{\|v_{j}\otimes v_{i}\|}.$
Therefore, given a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, there exists a unique polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ whose vertices satisfy (11). Let $l_{rs}$ denote the length of the ideal edge $\{rs\}\in E$, and let $\theta^{rs}_{t}$ denote the length of the boundary arc $t\in B$.
@@ 2.2. A geometric interpretation
The vectors $c_{ijk}$, $c_{rs}$, and $v^{\prime}_{t}$ lie in the same 2-dimensional time-like vector subspace of the Lorentzian space $\mathbb{R}^{3}$, where the notations $c_{ijk}$ and $c_{rs}$ are defined in Remark 1.4 and Remark 1.3, respectively.
@@
Without loss of generality, we assume $r=i,\ s=j,\ t=k$. Since $c_{ij}\in E_{ij}$, by Proposition 2.3, there exists a unique geodesic in $\mathbb{H}^{2}$ passing through $c_{ij}$ and Lorentz orthogonal to $E_{ij}$, which is $P_{ij}=\textrm{Span}(c_{ij},c_{ijk})\cap\mathbb{H}^{2}$. Note that $E_{ij}=(v^{\prime}_{k})^{\bot}\cap\mathbb{H}^{2}$. By Proposition 2.4, since $E_{ij}$ and $P_{ij}$ intersect at a right angle, it follows that $v^{\prime}_{k}\in\textrm{Span}(c_{ij},c_{ijk})$. Q.E.D.
Combining Remark 1.4 and Lemma 2.5, we obtain $c_{ijk}=\textrm{Span}(v^{\prime}_{i},c_{jk})\cap\textrm{Span}(v^{\prime}_{j},c_{ik})\cap\textrm{Span}(v^{\prime}_{k},c_{ij})$. Given the equivalence between $v_{i}v_{j}v_{k}$ and $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it is natural to consider a dual model. Note that for any $c_{ijk}\in\mathbb{R}^{3}$, the intersection $\textrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}$ may not lie within $\mathbb{H}^{2}$. In this paper, we focus on the case that
(12) $\textrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}\cap\mathbb{H}^{2}=\textrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}\neq\emptyset,\ \forall r\in\{i,j,k\}.$
The condition (12) constrains the position of $c_{ijk}$ when it is space-like, namely, $c_{ijk}$ can only lie within certain specific open domains. Let $c^{\prime}_{st}=\textrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}$ denote the dual edge center, which is the unique point on the geodesic $E^{\prime}_{st}$ such that the signed distance from $c^{\prime}_{st}$ to $\{rt\}$ is $\theta_{st}$, and the signed distance from $c^{\prime}_{st}$ to $\{rs\}$ is $\theta_{ts}$. Thus
@@
Note that the signed distance $\theta_{st}$ is positive if $c^{\prime}_{st}$ and $\{rs\}$ lie on the same side of $\{rt\}$ along the hyperbolic geodesic $E^{\prime}_{st}$, and negative otherwise. The above construction yields the following lemma.
@@
The vectors $c_{ijk}$, $c^{\prime}_{rs}$, and $v_{t}$ lie in the same 2-dimensional time-like vector subspace of $\mathbb{R}^{3}$.
By Proposition 2.3, let $P^{\prime}_{ij}$ denote the edge perpendicular passing through $c^{\prime}_{ij}$ and Lorentz orthogonal to $E^{\prime}_{ij}$. Then $P^{\prime}_{ij}=\textrm{Span}(v_{k},c_{ijk})\cap\mathbb{H}^{2}$. Consequently, the three edge perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ intersect in a common point $c_{ijk}$.
Recall the definition of the signed distance $h_{rs}$ from the face center $c_{ijk}$ to the hyperbolic geodesic $E_{rs}$. It is positive if $c_{ijk}$ lies on the same side of $E_{rs}$ as the right-angled hyperbolic hexagon $\{ijk\}$, negative otherwise, and zero if $c_{ijk}$ lies on $E_{rs}$. Similarly, let $q_{rs}$ denote the signed distance from $c_{ijk}$ to $E^{\prime}_{rs}$, which is defined to be positive if $c_{ijk}$ lies on the same side of the geodesic $E^{\prime}_{rs}$ as the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, negative otherwise, and zero if $c_{ijk}$ lies on $E^{\prime}_{rs}$. Please refer to Figure 3.
@@
In the Klein model, the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and its polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ partition the unit disk into 13 open domains, denoted by $D_{1},D_{2},...,D_{13}$ respectively. Please refer to Figure 4.
Figure 4. Partition of the right-angled hyperbolic hexagon $\{ijk\}$ in the Klein model. If $c_{ijk}$ is space-like, it may lie in the open domains bounded by two red tangents or two blue tangents.
By the definitions of $h$ and $q$, one can directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 13 open domains, as shown in Table 1. For example, if $c_{ijk}\in D_{1}$, then $c_{ijk}$ lies on the opposite side of $E_{ik}$ as $v_{i}v_{j}v_{k}$, so $h_{ik}<0$. Moreover, $c_{ijk}$ lies on the same side of $E_{ij}$ and $E_{jk}$ as $v_{i}v_{j}v_{k}$, which implies $h_{ij}>0$ and $h_{jk}>0$. Similarly, $c_{ijk}$ lies on the opposite side of $E^{\prime}_{ij}$, and the same side of $E^{\prime}_{ik}$ and $E^{\prime}_{jk}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, which implies $q_{ij}<0$, $q_{jk}>0$, and $q_{ik}>0$. If $c_{ijk}$ is space-like, it can lie only in one of 12 open domains, each bounded by two tangent lines and the unit circle. These domains are denoted by $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ and $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ respectively. Specifically, the domain $D_{\mathrm{i}}$ is enclosed by the tangents from vertices $v_{k}$ and $v_{j}^{\prime}$ to the unit circle, together with the unit circle itself. The domain $D_{\mathrm{I}}$ is enclosed by the tangents from vertices $v_{i}^{\prime}$ and $v_{j}^{\prime}$ to the unit circle, together with the unit circle itself. The remaining domains are defined analogously. Please refer to Figure 4. If $c_{ijk}$ does not lie in any of these 12 open domains, then either $\textrm{Span}(c_{ijk},v_{r})\cap\mathbb{H}^{2}=\emptyset$ or $\textrm{Span}(c_{ijk},v^{\prime}_{r})\cap\mathbb{H}^{2}=\emptyset$, which contradicts the assumption that the edge centers $c_{st}\in E_{rs}$ and $c^{\prime}_{st}\in E^{\prime}_{rs}$ are in $\mathbb{H}^{2}$ for $\{r,s,t\}=\{i,j,k\}$. Similar to the case that $c_{ijk}$ is time-like, one can also directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 12 open domains, as shown in Table 1. By continuously moving $c_{ijk}$ to the light cone, analogous results are obtained. Note that if $c_{ijk}$ is light-like, then it can lie only on the unit disk with 12 points removed.
@@
In the following, we present an alternative derivation of Table 1. Suppose $c_{ijk}$ is time-like. Without loss of generality, we assume $c_{ijk}\in D_{1}$. By Lemma 2.5, the vectors $c_{ijk}$, $c_{ik}$, and $v^{\prime}_{j}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Since $c_{ijk}\notin v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ik}\notin\{ik\}$. Furthermore, since $c_{ik}$ lies on the opposite side of $k$ as $i$. by the definition of $d$ in Remark 1.3, we have $d_{ki}<0$ and $d_{ik}>0$. Similarly, $d_{kj}<0$, $d_{jk}>0$, $d_{ij}>0$, and $d_{ji}>0$. By Lemma 2.6, the vectors $c_{ijk}$, $c^{\prime}_{ij}$, and $v_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Since $c_{ijk}\notin v_{i}v_{j}v_{k}$, it follows that $c^{\prime}_{ij}\notin k$. Furthermore, since $c^{\prime}_{ij}$ is on the opposite side of $\{ik\}$ as $\{jk\}$, it follows that $\theta_{ji}<0$ and $\theta_{ij}>0$. Similarly, $\theta_{jk}<0$, $\theta_{kj}>0$, $\theta_{ik}>0$, and $\theta_{ki}>0$. By Corollary 2.10, we have $h_{ik}<0$, $h_{jk}>0$, $h_{ij}>0$, $q_{ij}<0$, $q_{ik}>0$, and $q_{jk}>0$. The case that $c_{ijk}$ is space-like is established via the two lemmas below.
@@
The signs of $q$ and $h$ in the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ coincide with those in the domains $D_{2},D_{4},D_{6},D_{8},D_{10},D_{12}$, respectively.
@@
We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{V}}$ coincide with those in the domain $D_{10}$, and the remaining cases follow analogously.
If $c_{ijk}\in D_{10}$, then $q_{jk}<0$ and the rest are positive by Table 1. Suppose $c_{ijk}\in D_{\mathrm{V}}$. By Lemma 2.5, the vectors $c_{ijk}$, $c_{ij}$, and $v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Since $c_{ijk}\notin v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it follows that $c_{ij}\notin\{ij\}$. Furthermore, since $c_{ij}$ lies on the opposite side of $j$ as $i$, it follows that $d_{ij}<0$ and $d_{ji}>0$. Similarly, $d_{ik}<0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$. By Proposition 2.11, we have $h_{ij}>0$, $h_{jk}>0$, and $h_{ik}>0$. Note that $c_{ijk}\in v_{i}v_{j}v_{k}$. By Lemma 2.6, the vectors $c_{ijk}$, $c^{\prime}_{rs}$, and $v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$ for $\{r,s,t\}=\{i,j,k\}$. Consequently, $c^{\prime}_{rs}\in t$. Hence, $\theta_{ij}>0$, $\theta_{ji}>0$, $\theta_{jk}<0$, $\theta_{kj}>0$, $\theta_{ik}>0$, and $\theta_{ki}>0$. By Proposition 2.11, we have $q_{ij}>0$, $q_{jk}<0$, and $q_{ik}>0$. Q.E.D.
@@
The signs of $q$ and $h$ in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ coincide with those in the domains $D_{1},D_{3},D_{5},D_{7},D_{9},D_{11}$, respectively.